Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name jroche
Log entry time 16:19:44 on March 23, 2014
Entry number 387425
keyword=day shift summary
March 23, Day shift summary
Present:
J. Roche (SL)
S. Allison (3rd)
Guest apparence: K. Allada, M. Spata, Y. Roblin, D. Gaskell, B. Wojtsekhowski, L. Ou, D. Nguyen,
H. Fansler, A. Lumanog,...
Highlights:
- We got about 1.5 h of beam, that we used mostly to increase rates in the LHRS.
- Aborted attemp to move RHRS to small angle:
We started the day shift with an access to the hall in order to move the Right HRS from 33 deg to
15 deg. Once in the hall, Kalyan, Heidi and Andrew determined that moving the spectrometer
would requiere moving material inside a roped radiation area which would requiere radcon to
come in and take time. They decided to abort the mission.
- Modified magnet settings to improve rates: Changed central momentum of the magnets
from 2 GeV to 1 GeV. The LHRS went from ~25 Hz at 10 uA on C target to ~80 Hz in the same
condition, no effect on the RHRS. We also found that the sieve slits were in postion. We move the
small collimator in place for the LHRS but saw no improvment in the rates....
- Cryo event in the North Linac: around 12:30 the North Linac lost its cryo supply. We have
not received beam since. MCC projected around 1pm that beam will back around 5pm.
- Right around noon both HRS tripped : we did reset the quadrupoles by momentum and both
dipole by current. There has been no beam since the trip.
- Gaskell and Roblin were suposed to try to thread the beam in the Compton chicane but it did
not happen because of the cryo issue.
Issue requiring attention:
- No alarm when the HRSs tripped!!
Analysis: